What if you only know the 1st and 3rd value of a term value and you need to find the arithmetic explicit equation. How would you find the multiplier? What is the tactic to find that value without the guess and check method? Thank you very much. The table is down below.
__n__|_t(n)_
__1__|__17__
__2__|______
__3__|___3__
__4__|______

Answer by rothauserc(4718)

we are given two equations in 2 unknowns
:
1) a(1) = 17
:
a(n) = a(1) + d * (n - 1)
:
2) a(3) = a(1) + d * (3 - 1)
3 = 17 + 2d
2d = -14
d = -7
:
**********************
t(n) = 17 - 7 * (n -1)
:
t(2) = 17 - 7 = 10
:
t(4) = 17 - 21 = -4
**********************
: